OpenStudy (anonymous):

What is the slope of the line that contains the points (-6, 2) and (4, -5)?

OpenStudy (hoblos):

the slope between two points (x1,y1) and (x2,y2) is (y2-y1)/(x2-x1) so here the slope = (-5-2)/(4+6) = -7/10